OpenStudy (anonymous):

the chemistry lab at the university of Hardwoods keeps two acid solutions on hand. One is 20% acid and the other is 35% acid . how much 20% acid solution and how much 35% acid solution should be used to prepare 25 liters of a 26% acid solution?

OpenStudy (***[isuru]***):

hi, if we take a liter of 20% acid solution ... then that liter is 20% acid if we take a liter of 35% acid solution ... then that liter is 35% acid but wt the question ask is to prepare a solution of 26%acid with the volume of 25 liters. First we can calculate the amount we need for 1 liter.. and then we can multiply it by 25... if we take "x" amount of 20% the amount of 35% is (1-x) liters so.... 20x + 35(1 - x) = 26 20x + 35 -35x = 26 15x = 9 x = 9/15 liters x = 3/5 liters x = 600 ml so... to prepare a 1 liter solution of 26%acid u need 600 ml of 20% acid and the amount of 35% acid is 400 ml but to prepare a solution of 25 liters u will need ( 600ml x 25) of 20% acid which is = 15 liters and (400ml x 25) of 35% acid which is = 10 liters Hope this will help ya!!!!

OpenStudy (anonymous):

where did you get the 400 from

OpenStudy (***[isuru]***):

in the beginning I said that if the amount of 20% acid need for i liter is "X" then the amount of 35% acid that will be needed is (1-X) liters.... we got the value 3/5 for "X" so the amount of 35% is (1 - X) liters = (1 -3/5) liters = 2/5 liters which is 400 ml got it ?

OpenStudy (anonymous):

oh ok

OpenStudy (***[isuru]***):

Did u got the whole explanation now,,,, :-) ?

OpenStudy (anonymous):

yes